# Exploring Multiple Strategies to Improve Multilingual Coreference Resolution in CorefUD

### v2

@@ 7 Conclusion
We proposed several extensions to enhance the baseline models, including cross-lingual training, Span2Head modeling, syntactic information integration, headword mention representation, and long-context prediction. Among these, the heads-only model and singleton modeling showed the most consistent improvements across different datasets, demonstrating the importance of targeted adaptations for coreference resolution tasks. For several datasets, long-context prediction brings also a significant improvement. Additionally, our zero-shot cross-lingual experiments provided insights into the challenges and opportunities of cross-lingual transfer, with the Turkish dataset results highlighting potential issues related to noise in the training data.
Overall, our findings emphasize the need for tailored approaches in coreference resolution, particularly when dealing with diverse languages and annotation schemes. Future work could further investigate the impact of dataset-specific characteristics on model performance and explore additional strategies for enhancing cross-lingual transfer, especially for low-resource languages. Our source codes are publicly available for subsequent research888https://github.com/ondfa/coref-multiling.

### v3

@@ 7 Conclusion
We proposed several extensions to enhance the baseline models, including cross-lingual training, Span2Head modeling, syntactic information integration, headword mention representation, and long-context prediction. Among these, the heads-only model and singleton modeling showed the most consistent improvements across different datasets, demonstrating the importance of targeted adaptations for coreference resolution tasks. For several datasets, long-context prediction also brings a significant improvement. Additionally, our zero-shot cross-lingual experiments provided insights into the challenges and opportunities of cross-lingual transfer, with the Turkish dataset results highlighting potential issues related to noise in the training data.
Overall, our findings emphasize the need for tailored approaches in coreference resolution, particularly when dealing with diverse languages and annotation schemes. Future work could further investigate the impact of dataset-specific characteristics on model performance and explore additional strategies for enhancing cross-lingual transfer, especially for low-resource languages. Our source codes are publicly available for subsequent research999https://github.com/ondfa/coref-multiling.